OpenStudy (anonymous):

If (82)p = 84, what is the value of p?

OpenStudy (anonymous):

Divide both the sides by \(82\) to find \(p\)..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

No ok.. Just do it and show me what you did actually..

OpenStudy (anonymous):

2? dont you mean subtract?

OpenStudy (anonymous):

No, I mean divide..

OpenStudy (anonymous):

1.024.. if i divide

OpenStudy (anonymous):

If (8^2)p = 84, what is the value of p? nevermind this is the problem sorry

OpenStudy (anonymous):

i just looked at it

OpenStudy (anonymous):

Oh..

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

@NightCrawlers_

OpenStudy (anonymous):

Tell me : \(8^2 = ??\)

OpenStudy (nightcrawlers_):

She looks like she got this ^

OpenStudy (anonymous):

She? Who is she here?

OpenStudy (nightcrawlers_):

Oh he?

OpenStudy (anonymous):

ok back to the question

OpenStudy (anonymous):

Who is he here now?

OpenStudy (anonymous):

\[(8^{2})^{p}=84\]

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

Hey man..!! First write your question properly. Is that fine now? Is that now a right question?

OpenStudy (anonymous):

p= A2 B3 C4 D6

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

your question is still wrong..

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

It must be then: \[(8^2)^p = 8^4\]

OpenStudy (anonymous):

Otherwise, you won't get integer value for \(p\)..

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

but how would i work this out?

OpenStudy (anonymous):

See, when bases are same, exponents are equal.

OpenStudy (anonymous):

\[(a)^b = (a)^c \implies b = c\]

OpenStudy (anonymous):

p=2 then?

OpenStudy (anonymous):

Good.. Have you guessed it?

OpenStudy (anonymous):

yes sadly

OpenStudy (anonymous):

i wanna know how i did it or how to do it

OpenStudy (anonymous):

See, you need to just make the base look equal on both the sides..

OpenStudy (anonymous):

For that, look here, we have one power rule property: \[(a^b)^c = (a)^{b \times c}\]

OpenStudy (anonymous):

Now using it can I write : \[(8)^4 = (8)^{2 \times 2} \implies (8^2)^2\] See, this it is just a crystal clear use of that property..

OpenStudy (anonymous):

\[(8^{2})^{2}=8^{4} so then 8^{2} x ^{2}=8^{4}\]

OpenStudy (anonymous):

That is not \(x^2\), but there should be \(8^2\) instead of that..

OpenStudy (anonymous):

what i just wrote but better haha

OpenStudy (anonymous):

x=multiply

OpenStudy (anonymous):

p=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You can write it the best way as: \[(8^{2})^{2}=8^{4} \quad \text{so then} \quad 8^{2} \times 8 ^{2}=8^{4}\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\(\dagger\)